Paynesville Corporation manufactures and sells a preservative used in food and drug manufacturing. The company carries no inventories. The master budget calls for the company to manufacture and sell 116,000 liters at a budgeted price of $195 per liter this year. The standard direct cost sheet for one liter of the preservative follows.

Direct materials (2 pounds @ $12) $ 24
Direct labor (0.5 hours @ $40) 20


Variable overhead is applied based on direct labor hours. The variable overhead rate is $100 per direct-labor hour. The fixed overhead rate (at the master budget level of activity) is $18 per unit. All non-manufacturing costs are fixed and are budgeted at $2 million for the coming year.


At the end of the year, the costs analyst reported that the sales activity variance for the year was $606,000 unfavorable.

The following is the actual income statement (in thousands of dollars) for the year.

Sales revenue $ 21,718
Less variable costs
Direct materials 2,368
Direct labor 2,210
Variable overhead 5,230
Total variable costs $ 9,808
Contribution margin $ 11,910
Less fixed costs
Fixed manufacturing overhead 2,330
Non-manufacturing costs 1,310
Total fixed costs $ 3,640
Operating profit $ 8,270


Required:

What are the fixed overhead price and production volume variances for Paynesville? (Enter your answers in whole dollars. Indicate the effect of each variance by selecting "F" for favorable, or "U" for unfavorable. If there is no effect, do not select either option.)
Answer #1

sales activity variance = (actual sales volume - budgeted sales volume) x budgeted price

= - 606,000 = (actual sales volume - 116,000) x 195

= -606,000 / 195 = actual sales volume - 116,000

= -3,108 (rounded fig) = actual sales volume - 116,000

actual sales volume = 116,000 -3,108

actual sales volume = 112,892 (rounded fig)

Fixed overhead spending variance = actual fixed overhead - budgeted fixed overhead applied

= 3,640,000 - 4,032,056 = - 392,056 (F)

actual fixed overhead is the total fixed costs given.

budgeted fixed overhead cost = (standard rate x actual outputs) + non-manufacturing costs

= (18 x 112,892) + 2,000,000 = 2,032,056 + 2,000,000 = $ 4,032,056

fixed overhead spending variance is the difference between actual overhead incurred and the budgeted overhead for the actual output. The answer we got is negative, which means that it is a favorable variance. It is favorable since the actual amount spent is less than the amount that was thought to have incurred.

Fixed production volume variance = Budgeted fixed overhead - Standard fixed overhead applied

= 2,032,056 - 2,088,000 = - 55,944 (F)

budgeted fixed overhead = actual units x standard rate = 112,982 x 18 = 2,032,056

standard fixed overhead = standard units x standard rate = 116,000 x 18 = 2,088,000